OpenStudy (samigupta8):

For z1=[(1-i)/(1+i√3)]^(1/6) z2=[(1-i)/(√3+i)]^(1/6) z3=[(1+i)/(√3-1)]^(1/6) Which of the following holds? A).summation of |z|^2=3/2 B).summation of |z1|^3+|z2|^3=|z^3|^-6 C).|z1|^4+|z2|^4=|z3|^-8 D).|z1|^4+|z2|^4=|z3|^8

OpenStudy (samigupta8):

@ganeshie8

OpenStudy (samigupta8):

|Z3|^-6 in second option (rhs)

OpenStudy (samigupta8):

@ilovepuppieslol

OpenStudy (samigupta8):

@loser66

OpenStudy (loser66):

I can't read it . Can you take a snapshot?

OpenStudy (samigupta8):

I use os on mobile..

OpenStudy (samigupta8):

@snowsurf

OpenStudy (snowsurf):

This looks like a complex analysis problem. I'm not sure about this one.

OpenStudy (samigupta8):

@hartnn

hartnn (hartnn):

all the z's have magnitude 1/sqrt 2 isn't it ? also, whats the denominator of z3 ?

hartnn (hartnn):

i mean magnitude as \((\dfrac{1}{\sqrt 2})^{\dfrac{1}{6}}\)

OpenStudy (samigupta8):

Ty..